Please write an equation that satisfies the given conditions, and please show it step by step!!
Perpendicular to 3x+4y=12 and passes through (7,1)
Thank you :)

Answer by ewatrrr(24785)


Hi
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
Line Perpendicular to 3x+4y=12 and passes through (7,1)
3x+4y=12 OR y = (-3/4)x+3 ( m = (-3/4))
NEw Line: Perpendicular Line will have slope that is the negative reciprocal,
y = (4/3)x + b |Using ordered pair (7,1) to solve for b
1 = (4/3)*7 + b
-25/3 = b
y = (4/3)x - 25/3